Circular No. 5961

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994I IN NGC 5194 Several independent reports have been received of the discovery of

a supernova some 14′′ east and 12′′ south of the nucleus of NGC 5194 = M51: Apr. 2.17 UT, mag 13.5 (Tim Puckett and Jerry Armstrong, Atlanta, GA; CCD discovery; α = 13^{h}27^{m}47^{s}.62, δ = +47^{o}26′59′′.1, equinox 1950.0); 2.19, 13.7 (Wayne Johnson and Doug Millar, Anza, CA; visual discovery with CCD confirmation); 2.21, – (Richard Berry, Cedar Grove, WI; CCD discovery); 2.66, 13.8 (Reiki Kushida, Yatsugatake South Base Observatory; visual). M. W. Richmond and A. V. Filippenko have confirmed the object using BV RI CCD images obtained with the 0.5-m Berkeley Automatic Imaging Telescope at Leuschner Observatory and note it to be very blue. L. Armus and J. M. Mazzarella obtained moderate-dispersion spectra with the 5-m Hale reflector at Palomar Observatory and see broad undulations superposed on a generally featureless continuum.

PECULIAR VARIABLE IN SAGITTARIUS On Mar. 28 T. Hirayama, National Astronomical Observatory, Tokyo,

reported the discovery by M. Wakuda (Ryuyu, Shizuoka) on Mar. 14.825 UT of an object of mag ∼ 10.7 (T-Max 400 film, green filter) at α = 18^{h}51^{m}43^{s}, δ = −19^{o}45′.9 (equinox 1950.0, uncertainty 10′′). T. Kato, Kyoto University, reports that M. Yamamoto (Okazaki, Aichi) has photographic records showing the object’s apparent variability over mag 10.5– 11.5 (PO0 filter) since 1993 Mar. 29 but no earlier detection (and no brighter than mag 12.5 through 1992, according to Wakuda). Preliminary reduction of photometry by A. C. Gilmore at Mt. John Observatory on 1994 Mar. 30.72 UT gives V = 10.61, U −B = −0.23, B −V = +0.50, V −R = +0.35, V − I = +0.60 (uncertainty generally ±0.05 but up to ±0.1 in U − B).

E. K. Grebel, European Southern Observatory; and H. W. Duerbeck, Astronomical Institute, Münster, report spectroscopic observations on Mar. 30.4 with the ESO/MPI 2.2-m telescope and EFOSC-II. A low-resolution spectrum (range 340–920 nm, resolution 0.84 nm) shows emission lines of Hα, Hβ and the infrared Ca II (2) triplet, superimposed on a somewhat veiled late-K-giant continuum. A high-resolution spectrum (range 352–547 nm, resolution 0.19 nm) shows the Ca II H and K lines in absorption, as well as narrow (FWHM < ±250 km/s) emission lines of Hβ and Fe II lines (multiplets 41, 42, 48 and 49). The absence of blueshifted absorption lines and the narrowness of the emission lines suggest that the object is a symbiotic nova in a slightly progressed state.

1994 April 2 Brian G. Marsden